# The cohomological Kudla conjecture for unitary Shimura varieties

### v1

@@ 4.2. Completions of theta series of homogeneous polynomials
$f^{g}(\underline{\lambda}.A)=|\det(A)|^{2}f^{g}(\underline{\lambda})\leavevmode\nobreak\ .$
Therefore, we can apply the modularity results from Section 2.4 to $f^{g}$. Since $\Delta(f)=D$, then for $1\leq\ell\leq g$, we have:
$\Delta^{\ell}(f^{g})(\underline{\lambda})$ $=g!\ell!\sum_{\underset{|I|=g-\ell}{I\subseteq\{1,\ldots,g\}}}\bigwedge_{i\in I}f(\lambda_{i})\wedge D^{\ell}$
$=\frac{g!\ell!}{(g-\ell)!}\mathrm{Tr}\left(\Lambda^{g-\ell}((f(\lambda_{i},\lambda_{j}))_{1\leq i,j\leq g})\right)\wedge D^{\ell}$
where $\Delta$ is the Laplacian with respect to $\underline{\lambda}$, and for a $g\times g$ matrix $P$, $\Lambda^{\ell}(M)$ is the matrix of $(\ell)\times(\ell)$ minors. In particular,
$\Delta^{g}(f^{g})=(g!)^{2}D^{g}\leavevmode\nobreak\ .$
By Theorem 2.12, the generating series:
$\vartheta_{f^{g}}(\tau)=\det(Y)^{-1}\sum_{\underline{\lambda}\in(M^{\vee})^{g}}\exp(-\frac{\Delta}{4\pi})\left(f^{g}\right)(\underline{\lambda}\cdot Y^{\frac{1}{2}})q^{Q(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\in{\mathcal{H}}^{1,1}(E_{M},\mathbb{R})[(M^{\vee}/M)^{g}]\llbracket q\rrbracket\leavevmode\nobreak\ ,$
@@
$g!\leavevmode\nobreak\ \sum_{\underline{\lambda}\in(M^{\vee})^{g}}\varphi(\tau,\underline{\lambda})q^{h(\underline{\lambda})}\leavevmode\nobreak\ .$
@@
$\varphi(\tau,\underline{\lambda})$ $=\frac{1}{\det(Y)}\sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\Bigg{(}\Lambda^{g-\ell}\Big{(}\sqrt{Y}[f(\lambda_{i},\lambda_{j})]_{1\leq i,j\leq g}\sqrt{Y}\Big{)}\Bigg{)}\wedge D^{\ell}$
$=\leavevmode\nobreak\ \sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\Big{(}\Lambda^{\ell}(Y^{-1})\Lambda^{g-\ell}([f(\lambda_{i},\lambda_{j})]_{1\leq i,j\leq g})\Big{)}\wedge D^{\ell}$
@@
$[\varphi(\tau,\underline{\lambda})]=\sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\left(\Lambda^{\ell}(Y^{-1})\Lambda^{g-\ell}([(f(\lambda_{i},\lambda_{j}))_{1\leq i,j\leq g}]\right)\wedge D^{\ell}\leavevmode\nobreak\ .$
@@
$\frac{1}{(g-\ell)!}\Lambda^{g-\ell}\Big{(}(f(\lambda_{i},\lambda_{j}))_{1\leq i,j\leq g}\Big{)}$
@@
The generating series $\Phi_{M}^{g}$ admits the non-holomorphic completion:
@@
that transforms like a Hermitian modular form of weight $2+n$ with respect to the Weil representation $\rho$. In particular, it is a Hermitian quasi-modular form of weight $1+n$ with respect to the Weil representation $\rho_{M}$.
@@
$-\frac{1}{4\pi y}\Theta_{M}\wedge D\leavevmode\nobreak\ ,$
@@
Let $[\widetilde{Z}(\lambda)]=[Z(\lambda)]-\frac{h(\lambda)}{n}D$. Then the generating series
@@ 4.3. Correction to the theta series
Corollary 4.13 constructs a natural holomorphic correction of special divisors that provide holomorphic modular forms and our goal in this section is to provide similar holomorphic corrections in higher codimension, which are not afforded by Theorem 4.12.
@@
$\mathrm{H}^{g,g}(E_{M},\mathbb{Q})\simeq\oplus_{\ell\leq g}{\mathcal{L}}^{\ell}\mathrm{H}_{\mathrm{prim}}^{g-\ell,g-\ell}(E_{M},\mathbb{Q})\leavevmode\nobreak\ .$
@@
$\underline{\lambda}$ $\longmapsto Z(\underline{\lambda})\cdot\alpha=\int_{\alpha}f(\lambda_{1})\wedge\cdots\wedge f(\lambda_{g})\leavevmode\nobreak\ .$
By Proposition 4.9, the above polynomial is an element of ${\mathcal{F}}_{n,g}$. In other words, we get an element in $u_{g}\in\mathrm{H}^{g,g}(E_{M},\mathbb{R})\otimes{\mathcal{F}}_{n,g}$. Putting together these maps for $0\leq g\leq n$, we get an element in
$\bigoplus_{g=0}^{n}\mathrm{H}^{g,g}(E_{M},\mathbb{R})\otimes{\mathcal{F}}_{n,g}\leavevmode\nobreak\ .$
@@
The linear map $u:\bigoplus_{g=0}^{n}\mathrm{H}_{g,g}\rightarrow\bigoplus_{g=0}^{n}{\mathcal{F}}_{n,g}$ is a morphism of $\mathfrak{sl}_{2}$ representations. In other words, $\mathfrak{sl}_{2}$ acts trivially on $u$.
@@
Indeed, let $\alpha\in\mathrm{H}_{g,g}(E_{M},\mathbb{Z})$. Then ${\mathcal{L}}^{*}(\alpha)$ by definition acts on $\mathrm{H}^{g-1,g-1}$ as the linear form $\beta\mapsto\beta\cup D\cup\alpha$. Therefore, for any $\lambda_{1},\lambda_{g-1}$, we have $u_{g-1}(\alpha)(\lambda_{1},\ldots,\lambda_{g-1})=\alpha\cup D\cup Z(\lambda_{1},\ldots,\lambda_{g-1})$.
@@
$\Delta([Z(\lambda_{1},\ldots,\lambda_{g})])=[Z(\lambda_{1},\ldots,\lambda_{g-1})]\cup D,$
@@
$\Delta(u_{g})(\lambda_{1},\ldots,\lambda_{g-1})=\alpha\cup D\cup[Z(\lambda_{1},\ldots,\lambda_{g-1})],$
@@
Next, notice that $H$ acts both on ${\mathcal{F}}_{n,g}$ and $\mathrm{H}_{g,g}(E_{M},\mathbb{Z})$ by multiplication by $n-2g$. Therefore we conclude that $u$ is a morphism of $\mathfrak{sl}_{2}$ representations, since the equivariance under the third operator is automatic; see the corollary hereafter. ∎
@@
Since a morphism of $\mathfrak{sl}_{2}$ representations preserve the Lefschetz decompositions, we get the following corollary.
@@
$\bigoplus_{\ell=0}^{g}{\mathcal{L}}^{g-\ell}\mathrm{H}^{\ell,\ell}_{\mathrm{prim}}\otimes\Lambda^{g-\ell}{\mathcal{F}}_{n,\ell}^{\mathrm{prim}}\leavevmode\nobreak\ .$
For each $0\leq\ell\leq\frac{n}{2}$, let $(W^{\ell}_{i})_{i\in I_{\ell}}$ denote a basis of $\mathrm{H}^{\ell,\ell}_{\mathrm{prim}}(E_{M},\mathbb{Q})$. For $\ell=0$, $\mathrm{H}^{0,0}_{\mathrm{prim}}(E_{M},\mathbb{Q})$ is one dimensional and we will take $W^{0}$ to be simply the fundamental class of $E_{M}$. By Corollary 4.16, we can decompose $[Z(\lambda_{1},\ldots\lambda_{g})]$ as follows:
$[Z(\lambda_{1},\ldots,\lambda_{g})]=\sum_{\ell=0}^{g}\sum_{i\in I_{\ell}}P^{\ell,g}_{i}(\lambda_{1},\ldots,\lambda_{g})W^{\ell}_{i}\otimes D^{g-\ell}$
where $P_{i}^{g,\ell}(\lambda_{1},\ldots,\lambda_{g})\in\Lambda^{g-\ell}{\mathcal{F}}_{n,\ell}^{\mathrm{prim}}$. In particular, the decomposition along ${\mathcal{F}}^{\mathrm{prim}}_{n,g}$ is of the form: $\sum_{i\in I_{g}}P_{i}^{g,g}(\lambda_{1},\ldots,\lambda_{g})W^{g}_{i}$ where the polynomials $P_{i}^{g,g}(\lambda_{1},\ldots,\lambda_{g})\in{\mathcal{F}}^{\mathrm{prim}}_{n,g}$ are harmonic polynomials.
@@
    For each $0\leq\ell\leq\frac{g}{2}$, and $i\in I_{\ell}$ $P^{\ell}_{i}$ is a harmonic polynomial on $(M_{\mathbb{R}})^{\ell}$.
@@
    The following relations holds:
    $\Delta_{g}P_{i}^{g,\ell}(\lambda_{1},\ldots,\lambda_{g})=P_{i}^{g,\ell}(\lambda_{1},\ldots,\lambda_{g-1})\leavevmode\nobreak\ .$
    and
    $\Lambda P_{i}^{g-1,\ell}=P^{g,\ell}_{i}\leavevmode\nobreak\ .$
@@
(1) follows from the Lefschtez decomposition and (2) follows from the relation:
$\Delta_{g}[Z(\lambda_{1},\ldots,\lambda_{g})]=[Z(\lambda_{1},\ldots,\lambda_{g-1})]\otimes D\leavevmode\nobreak\ ,$
and the second relation follows from Corollary 4.15. ∎
@@
We have $P^{g,\ell}_{i}=\Lambda^{g-k}P^{\ell,\ell}_{i}$ and $P^{\ell,\ell}_{i}$ is a pluriharmonic polynomial.
Let
$[\widetilde{Z}(\underline{\lambda})]=[Z(\underline{\lambda})]-\sum_{\ell=0}^{g-1}\sum_{i\in I_{\ell}}[\Lambda^{g-\ell}(P_{i}^{\ell,\ell})](\lambda_{1},\ldots,\lambda_{g})[W_{i}^{\ell}\cup D^{g-\ell}].$
@@
The generating series
$\sum_{\underline{\lambda}\in(M^{\vee})^{g}}[\widetilde{Z}(\underline{\lambda})]q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}=\Phi_{M}^{g}-\sum_{\ell=0}^{g-1}\sum_{i\in I_{\ell}}\vartheta_{\Lambda^{g-\ell}P^{\ell,\ell}_{i}}\leavevmode\nobreak\ [W^{\ell}_{i}\cup D^{g-\ell}]$
is a Hermitian modular form of weight $2+n$ with respect to the Weil representation $\rho_{L}$ of $U(n,n)(\mathbb{Z})$

### v2

@@ 4.2. Completions of theta series of homogeneous polynomials
$f^{g}(\underline{\lambda}.A)=|\det(A)|^{2}f^{g}(\underline{\lambda})\penalty 10000\ .$
Therefore, we can apply the modularity results from Section˜2.5 to $f^{g}$. Since $\Delta(f)=D_{\mathfrak{J}}$, then for $1\leq\ell\leq g$, we have:
$\Delta^{\ell}(f^{g})(\underline{\lambda})$ $=g!\ell!\sum_{\underset{|I|=g-\ell}{I\subseteq\{1,\ldots,g\}}}\bigwedge_{i\in I}f(\lambda_{i})\wedge D_{{\mathfrak{J}}}^{\ell}$
$=\frac{g!\ell!}{(g-\ell)!}\mathrm{Tr}\left(\wedge^{g-\ell}((f(\lambda_{i},\lambda_{j}))_{1\leq i,j\leq g})\right)\wedge D_{{\mathfrak{J}}}^{\ell}$
where $\Delta$ is the Laplacian with respect to $\underline{\lambda}$, and for a $g\times g$ matrix $P$, $\wedge^{\ell}(M)$ is the matrix of $\ell\times\ell$ minors. In particular,
$\Delta^{g}(f^{g})=(g!)^{2}D_{{\mathfrak{J}}}^{g}\penalty 10000\ .$
By Theorem˜2.12, the generating series:
$\vartheta_{f^{g}}(\tau)=\det(Y)^{-1}\sum_{\underline{\lambda}\in(M^{\vee})^{g}}\exp(-\frac{\Delta}{4\pi})\left(f^{g}\right)(\underline{\lambda}\cdot Y^{\frac{1}{2}})q^{Q(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}\in{\mathcal{H}}^{g,g}(E_{M},\mathbb{R})[(M^{\vee}/M)^{g}]\llbracket q\rrbracket\penalty 10000\ ,$
@@
$g!\penalty 10000\ \sum_{\underline{\lambda}\in(M^{\vee})^{g}}\varphi(\tau,\underline{\lambda})q^{h(\underline{\lambda})}\penalty 10000\ .$
@@
$\varphi(\tau,\underline{\lambda})$ $=\frac{1}{\det(Y)}\sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\Bigg(\wedge^{g-\ell}\Big(\sqrt{Y}[f(\lambda_{i},\lambda_{j})]_{1\leq i,j\leq g}\sqrt{Y}\Big)\Bigg)\wedge D_{{\mathfrak{J}}}^{\ell}$
$=\penalty 10000\ \sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\Big(\wedge^{\ell}(Y^{-1})\wedge^{g-\ell}([f(\lambda_{i},\lambda_{j})]_{1\leq i,j\leq g})\Big)\wedge D_{{\mathfrak{J}}}^{\ell}$
@@
$[\varphi(\tau,\underline{\lambda})]=\sum_{\ell=1}^{g}\frac{(-1)^{\ell}}{(4\pi)^{\ell}(g-\ell)!}\mathrm{Tr}\left(\wedge^{\ell}(Y^{-1})\wedge^{g-\ell}([(f(\lambda_{i},\lambda_{j}))_{1\leq i,j\leq g}]\right)\wedge D_{{\mathfrak{J}}}^{\ell}\penalty 10000\ .$
@@
$\frac{1}{(g-\ell)!}\wedge^{g-\ell}\Big((f(\lambda_{i},\lambda_{j}))_{1\leq i,j\leq g}\Big)$
@@
The generating series $\Phi_{M}^{g}$ admits the non-holomorphic completion in cohomology:
@@
that transforms like a Hermitian modular form of weight $2+n$ with respect to the Weil representation $\rho$. In particular, it is a Hermitian quasi-modular form of weight $2+n$ with respect to the Weil representation $\rho_{M}$.
@@
$\frac{1}{4\pi y}\Theta_{M}\wedge D_{{\mathfrak{J}}}\penalty 10000\ ,$
@@
Let $\widetilde{Z}(\lambda)]=[Z(\lambda)]-\frac{h(\lambda)}{n}D_{\mathfrak{J}}$. Then the generating series
@@ 4.3. Correction to the theta series
Corollary˜4.13 constructs a natural holomorphic correction of special divisors that provide holomorphic modular forms and our goal in this section is to provide similar holomorphic corrections in higher codimension, which are not afforded by Theorem˜4.12.
@@
$\mathrm{H}^{g,g}(E_{M},\mathbb{Q})\simeq\oplus_{\ell\leq g}{\mathcal{L}}^{\ell}\mathrm{H}_{\mathrm{prim}}^{g-\ell,g-\ell}(E_{M},\mathbb{Q})\penalty 10000\ .$
@@
$\underline{\lambda}$ $\longmapsto Z(\underline{\lambda})\cdot\alpha=\int_{\alpha}f(\lambda_{1})\wedge\cdots\wedge f(\lambda_{g})\penalty 10000\ .$
By Proposition˜4.9, the above polynomial is an element of ${\mathcal{F}}_{n,g}$. In other words, we get an element in $u_{g}\in\mathrm{H}^{g,g}(E_{M},\mathbb{R})\otimes{\mathcal{F}}_{n,g}$. Putting together these maps for $0\leq g\leq n$, we get an element in
$\bigoplus_{g=0}^{n}\mathrm{H}^{g,g}(E_{M},\mathbb{R})\otimes{\mathcal{F}}_{n,g}\penalty 10000\ .$
@@
The linear map $u:\bigoplus_{g=0}^{n}\mathrm{H}_{g,g}(E_{M},\mathbb{R})\rightarrow\bigoplus_{g=0}^{n}{\mathcal{F}}_{n,g}$ is a morphism of $\mathfrak{sl}_{2}$ representations. In other words, $\mathfrak{sl}_{2}$ acts trivially on $u$.
@@
${\mathrm{H}_{g,g}(E_{M},\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u_{g}}\scriptstyle{{\mathcal{L}}^{*}}$${{\mathcal{F}}_{n,g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Delta}$${\mathrm{H}_{g-1,g-1}(E_{M},\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u_{g-1}}$${{\mathcal{F}}_{n,g-1},}$
Indeed, let $\alpha\in\mathrm{H}_{g,g}(E_{M},\mathbb{Z})$. Then ${\mathcal{L}}^{*}(\alpha)$ by definition acts on $\mathrm{H}^{g-1,g-1}$ as the linear form $\beta\mapsto\beta\cup D_{\mathfrak{J}}\cup\alpha$. Therefore, for any $\lambda_{1},\lambda_{g-1}$, we have $u_{g-1}(\alpha)(\lambda_{1},\ldots,\lambda_{g-1})=\alpha\cup D_{{\mathfrak{J}}}\cup Z(\lambda_{1},\ldots,\lambda_{g-1})$.
@@
$\Delta([Z(\lambda_{1},\ldots,\lambda_{g})])=[Z(\lambda_{1},\ldots,\lambda_{g-1})]\cup D_{{\mathfrak{J}}},$
@@
$\Delta(u_{g})(\lambda_{1},\ldots,\lambda_{g-1})=\alpha\cup D_{{\mathfrak{J}}}\cup[Z(\lambda_{1},\ldots,\lambda_{g-1})],$
@@
Next, notice that $H$ acts both on ${\mathcal{F}}_{n,g}$ and $\mathrm{H}_{g,g}(E_{M},\mathbb{Z})$ by multiplication by n2gn-2g. Therefore we conclude that $u$ is a morphism of $\mathfrak{sl}_{2}$ representations, since the equivariance under the third operator is automatic; see the corollary hereafter. ∎
@@
${\mathrm{H}_{g,g}(E_{M},\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u_{g}}$$\scriptstyle{\delta^{*}}$${{\mathcal{F}}_{n,g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Lambda}$${\mathrm{H}_{g+1,g+1}(E_{M},\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u_{g+1}}$${{\mathcal{F}}_{n,g+1},}$
@@
Any morphism of $\mathfrak{sl}_{2}$ representations will preserve the Lefschetz decompositions, so we obtain the following corollary.
@@
$\bigoplus_{\ell=0}^{g}{\mathcal{L}}^{g-\ell}\mathrm{H}^{\ell,\ell}_{\mathrm{prim}}(E_{M},\mathbb{Q})\otimes\Lambda^{g-\ell}{\mathcal{F}}_{n,\ell}^{\mathrm{prim}}\penalty 10000\ .$
For each $0\leq\ell\leq\frac{n}{2}$, let $(W^{\ell}_{i})_{i\in I_{\ell}}$ denote a basis of $\mathrm{H}^{\ell,\ell}_{\mathrm{prim}}(E_{M},\mathbb{Q})$. For $\ell=0$, $\mathrm{H}^{0,0}_{\mathrm{prim}}(E_{M},\mathbb{Q})$ is 1-dimensional, and we take $W^{0}$ to be the fundamental class of $E_{M}$. By Corollary˜4.16, we can decompose $Z(\lambda_{1},\ldots\lambda_{g})]$ as follows:
$Z(\lambda_{1},\ldots,\lambda_{g})]=\sum_{\ell=0}^{g}\sum_{i\in I_{\ell}}P^{g,\ell}_{i}(\lambda_{1},\ldots,\lambda_{g})W^{\ell}_{i}\otimes D_{\mathfrak{J}}^{g-\ell}$
where $P_{i}^{g,\ell}(\lambda_{1},\ldots,\lambda_{g})\in\Lambda^{g-\ell}{\mathcal{F}}_{n,\ell}^{\mathrm{prim}}$. In particular, the decomposition along ${\mathcal{F}}^{\mathrm{prim}}_{n,g}$ is of the form: $\sum_{i\in I_{g}}P_{i}^{g,g}(\lambda_{1},\ldots,\lambda_{g})W^{g}_{i}$ where the polynomials
$P_{i}^{g,g}(\lambda_{1},\ldots,\lambda_{g})\in{\mathcal{F}}^{\mathrm{prim}}_{n,g}$
are harmonic polynomials.
@@
    For each $0\leq\ell\leq\frac{g}{2}$ and $i\in I_{\ell}$, $P^{\ell,\ell}_{i}$ is a harmonic polynomial on $(M_{\mathbb{R}})^{\ell}$.
@@
    The following relations hold:
    $\Delta_{g}P_{i}^{g,\ell}(\lambda_{1},\dots,\lambda_{g})$ $=P_{i}^{g,\ell}(\lambda_{1},\dots,\lambda_{g-1})\penalty 10000\ ;$
    $\Lambda P_{i}^{g-1,\ell}$ $=P_{i}^{g,\ell}\penalty 10000\ .$
@@
(1) follows from the Lefschetz decomposition. For (2), the first follows from the relation:
$\Delta_{g}[Z(\lambda_{1},\ldots,\lambda_{g})]=[Z(\lambda_{1},\ldots,\lambda_{g-1})]\cup D_{\mathfrak{J}}\penalty 10000\ ,$
and the second follows from Corollary˜4.15. ∎
@@
We have $P^{g,\ell}_{i}=\Lambda^{g-\ell}P^{\ell,\ell}_{i},$ and $P^{\ell,\ell}_{i}$ is a pluriharmonic polynomial.
Now, let
$\widetilde{Z}(\underline{\lambda})]=[Z(\underline{\lambda})]-\sum_{\ell=0}^{g-1}\sum_{i\in I_{\ell}}[\Lambda^{g-\ell}(P_{i}^{\ell,\ell})](\lambda_{1},\ldots,\lambda_{g})[W_{i}^{\ell}\cup D_{{\mathfrak{J}}}^{g-\ell}]\penalty 10000\ .$
@@
The generating series of cohomology classes
$\sum_{\underline{\lambda}\in(M^{\vee})^{g}}[\widetilde{Z}(\underline{\lambda})]q^{h(\underline{\lambda})}\mathfrak{e}_{\underline{\lambda}}=\Phi_{M}^{g}-\sum_{\ell=0}^{g-1}\sum_{i\in I_{\ell}}\vartheta_{\Lambda^{g-\ell}P^{\ell,\ell}_{i}}\penalty 10000\ [W^{\ell}_{i}\cup D_{{\mathfrak{J}}}^{g-\ell}]$
is a Hermitian modular form of weight $2+n$ with respect to the Weil representation $\rho_{L}$ of $U(g,g)(\mathbb{Z})$ .